OpenStudy (abbie):

help? solve for x. -3x+8y=5 6x-2y=10

OpenStudy (abbie):

A. x= 5/2 B. x=15/7 C. x=10/7 D. x=5/7

OpenStudy (abbie):

@Reaper534 @KamiBug

OpenStudy (lukebluefive):

Looks like the substitution method should work. Try getting "y" on one side by itself by adding terms to or subtracting terms from both sides.

OpenStudy (lukebluefive):

I'll get you started. Take the first equation: \[-3x+8y =5\] Add 3x to both sides: \[8y =5 + 3x\] Then divide both sides by 8: \[y =\frac{5 + 3x}{8}\] Now, plug this value for y into the second equation and solve for x.

OpenStudy (abbie):

which value?

OpenStudy (abbie):

@agent0smith

OpenStudy (lukebluefive):

Plug (5 + 3x)/8 in where y occurs in the second equation (6x - 2y = 10).

OpenStudy (abbie):

so 10= (5+3x)/8

OpenStudy (abbie):

wouls A be correct?

OpenStudy (agent0smith):

-3x+8y=5 6x-2y=10 easiest way is to to use elimination... multiply the second equation by 4 and add them together

OpenStudy (abbie):

how do i multiply it?

OpenStudy (agent0smith):

multiply everything by 4

OpenStudy (abbie):

I just guessed because my time was over I put C as an answer :/

OpenStudy (kamibug):

sorry i wasn't online! =(

OpenStudy (abbie):

its ok!